OpenStudy (anonymous):

find the ranger of values of m such as the straight line y = 2x + m intersects the curve x^2 + y^2 = 4 at two distinct points

OpenStudy (anonymous):

typo there range*

OpenStudy (dumbcow):

x^2 + y^2 = 4 is a circle with radius 2, centered at the origin  find the y-intercepts at those 2 tangent points to get range of m dy/dx = slope of tangent line = 2

OpenStudy (callisto):

Put y = 2x + m into x^2 + y^2 = 4 x^2 + (2x + m)^2 = 4 5x^2 + 4mx +m^2-4 =0 For it to have 2 distinct intersection points, Δ>0 (4m)^2 - 4(5)(m^2-4) >0 -4m^2 +80>0 4m^2<80 m^2 < 20 -sqrt20<m<sqrt20